### v1

# Distance Recomputator and Topology Reconstructor for Graph Neural Networks

Abstract
Graph Neural Networks (GNNs) have gained prominence in semi-supervised learning for graph representation due to their ability to capture intricate node relationships. Recently, there is a trend for k-hop structure learning for GNNs. While GAMLP [ZYS+22] trains an MLP layer for each k-hop domain, ImprovingTE [YWZL23] enhances this approach by injecting contextualized substructure information to effectively utilize the k-hop structure. However, those traditional k-hop sampling approaches have largely relied on sampling performance, which limits the upper bound of accuracy and made the outcome unstable. To address this limitation, inspired by ”coraset selection”[GZB22] idea, we develop a novel approach that facilitates k-hop structure sampling and message passing, extending the reach and depth of information flow within the graph. To tackle the challenges of mislabeling and inaccuracies in datasets, we introduce two innovative models: the ”Distance Recomputator” and the ”Topology Reconstructor.” The Distance Recomputator recalibrates the distances between nodes, thereby refining node representations and interactions in a more accurate and context-aware manner. Complementing this, the Topology Reconstructor dynamically adjusts local graph structures, enhancing the model’s adaptability to complex and evolving graph topologies. Our experimental results indicate significant performance enhancements of these models over existing benchmarks.
1University of Wisconsin-Madison,
2University of Notre Dame
1 Introduction
Graph Neural Networks (GNNs) have emerged as a powerful tool in the realm of machine learning, adept at capturing the complex relationships inherent in graph-structured data. From social network analysis to molecular structure interpretation, GNNs have demonstrated remarkable versatility. However, their ability to model dependencies and interactions within graph structures is fundamentally constrained by the methods used to compute and represent node relationships. Besides, GNN in graph local structure learning (K-hop) has became more and more popular, such as [WYHL21]. Traditional GNN architectures, while effective, often struggle with efficiently encoding the dynamic and intricate topologies of real-world graphs. This limitation motivates the need for more adaptive and robust models that can better capture the nuances of graph data.
Current GNN models predominantly rely on static node representations and fixed neighborhood aggregation schemes, leading to suboptimal performance in scenarios where graph topology is not only complex but also dynamic. Additionally, the computation of node distances in large graphs often incurs significant computational overhead, thereby limiting scalability. These challenges are accentuated in applications involving large-scale graphs or rapidly evolving network structures, such as in communication networks or dynamic social graphs. The limitations in handling varying node distances and adapting to topological changes prompt the exploration of more flexible and efficient approaches.
To address these challenges, we introduce two novel models: the ”Distance Recomputator” and the ”Topology Reconstructor.” The Distance Recomputator is designed to efficiently recalibrate node distances within a specified k-hop domain, leveraging a dynamic encoding scheme that adapts to changes in node proximity and graph density. This allows for a more nuanced representation of node relationships, enhancing the accuracy of dependency modeling in the network. Complementing this, the Topology Reconstructor dynamically adjusts the local network topology, enabling the model to respond to structural changes in the graph. By integrating these models into standard GNN frameworks, we propose a solution that not only addresses the static nature of traditional models but also introduces a level of adaptability hitherto unseen in GNN architectures.
Our experimental evaluation of these models demonstrates a marked improvement in performance across a range of benchmark datasets, particularly in tasks involving dynamic or large-scale graphs. The Distance Recomputator shows enhanced efficiency in recalculating node distances, leading to more accurate node representations and predictions. Similarly, the Topology Reconstructor proves effective in adapting to topological changes, thereby maintaining model robustness. Complementing these empirical results, we provide a comprehensive theoretical analysis that elucidates the mechanisms by which these models achieve superior performance. This analysis not only validates our experimental findings but also contributes to a deeper understanding of the underlying principles governing effective GNN design.
2 Motivation
2.1 Topology imbalance Problem
”Not every link are useful as their topology reflected”. The topology of a graph in GNNs does not always accurately reflect the significance of each connection. [LLC+23] Consider a social network: an account labeled ”math” might follow one food-maker and ten mathematical educators. In this context, the link to the food-maker is less relevant for label prediction tasks. Current graph neural models, such as GNNs, indiscriminately propagate, aggregate, and update information from all 1-hop neighbors, including less relevant connections like the food-maker. This issue highlights the ’Topology Imbalance Problem,’ where certain links (e.g., those to math educators) are more informative and relevant than others (e.g., the food-maker link).
Although models like Graph Attention Networks (GAT) assign an importance score to each link, they rely on a synchronous aggregator, treating all nodes within the same hop identically during propagation and update phases. To address this imbalance, our paper introduces the ’Distance Recomputator’ model, which recalculates node distances to better reflect the relevance of links. Complementarily, we propose an ’Asynchronous Aggregator’ that enables nodes to be aggregated based on these recalculated distances, allowing for more nuanced and context-sensitive information processing.
2.2 The Shortcoming of Synchronous Aggregator
Synchronous aggregators in Graph Neural Network process neighbors in one aggregation, the synchronous approach can inadvertently amplify the impact of less relevant or erroneous links, potentially degrading the model’s performance. It fails to differentiate between the varying levels of relevance among neighboring nodes, treating all connections as equally significant during the aggregation process. Our proposed asynchronous aggregator aims to mitigate this shortcoming by allowing for selective, relevance-based aggregation of neighborhood information, thereby enhancing the model’s accuracy and robustness against irrelevant or misleading connections.
Some current work shows the asynchronous processing of GNNs, such as AEGNN[SGS22], which designs update rules that restrict recomputation of network activations only to the nodes based on each new event, while Gated Graph Sequence Neural Networks [LTBZ17] deploys a method that can use gated recurrent units to extend to output sequences of GNNs to realize asynchronous processing. While all above methods give solution for asynchronous processing in message passing, they did not give the solution at a topology level.
2.3 Heat Diffusion in Graph Neural Networks
Heat Diffusion in GNNs is inspired by the physical phenomenon of heat transimisision, where heat (information) dissipates as it moves away from the heat source.[TDKF16] Our work also explores the concept of heat diffusion in the context of GNNs, particularly in k-hop message passing. We simulate thermodynamic properties, where information intensity decreases progressively during propagation. By applying this concept to GNNs, we introduce a novel method to control information flow, ensuring that the propagation of data mimics natural attenuation over distance. This method not only provides a more realistic approach to information dissemination in networks but also helps in reducing the noise and enhancing the signal-to-noise ratio in the message-passing process. More details about this innovative approach can be found in our supplementary document, ”Heat Diffusion in GNNs”.
More details in this link: Heat Diffusion in GNNs
2.4 Paper Contribution
  • K-hop Diffusion Message Passing Strategy: We develop a message passing framework enabling vertices to learn k-hop information at each propagation stage. Post-propagation, node representations and inter-node distances are recalibrated, reflecting the acquired k-hop information. This layer-by-layer information dissemination, incorporating heat diffusion within the k-hop domain, refines node characteristics and relational metrics continuously.
  • Distance Recomputator: Our model employs an attention mechanism to recalculate node distances, considering both k-hop topology and vertex features. This recalibration allows for accurate, context-aware node relationship representations within the graph.
  • Topology Reconstructor: Introducing the first k-hop topology reconstruction method, our model leverages k-hop topology from sampling to perform reconstruction based on computed ”similarity distances.” Nodes exceeding a similarity distance threshold are repositioned to optimize network configurations for enhanced learning outcomes.
3 Related Work and Innovation
3.1 K-hop Message Passing
The prevailing K-Hop models in graph neural networks predominantly emphasize K-time message passing with 1-hop propagation schemes. Existing models lack the capability for k-hop message passing with k-hop propagation in a single iteration, or for one-time k-hop sampling methods. For instance, the enhancing multi-hop connectivity model reevaluates connectivity in the k-hop neighborhood by sampling k-hop neighbors and reweighting high-order connections, rewarding highly-related nodes and penalizing less-correlated ones [LJZ+22]. Another model, KP-GNN, incorporates peripheral embeddings to enrich representation learning at each layer [FCL+22]. Our model breaks new ground by efficiently sampling k-hop neighborhood information and realizing k-hop message passing within this domain. By exploring k-hop neighborhoods in each propagation cycle, our framework captures more locality information than conventional 1-hop methods, thereby enhancing label prediction accuracy for the central node with a comprehensive k-hop perspective.
3.2 Diffusion Propagation in Graph Neural Networks
[…]
3.3 Graph Imbalancing Problems
The issue of label and topology imbalance in graph neural networks has been relatively underexplored. Zhao et al. [ZLZW22] proposed adjusting edge weights to mitigate topology imbalance effects without altering the graph structure. In contrast, our research introduces two innovative models to address this challenge comprehensively. The first, GKHDDRA, utilizes hop jumping to adjust the graph’s topology. The second, GDRA, fine-tunes the dataset by discarding irrelevant edges and forming new connections between strongly similar nodes, reshaping the graph’s topology to achieve a more balanced and representative structure.
4 Basic Methodology on Graph Neural Networks
Consider an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with the node set $\mathcal{V}$ and edge set $\mathcal{E}$. The adjancy matrix A $\in$ $R^{N\times N}$ describes the existence of connections among nodes. Every node $v_{i}\in\mathcal{V}$ has an associated feature vector $x_{i}\in R^{d\times 1}$, so the whole feature vector space can be represented as $X=[x_{1},x_{2},...,x_{N}]^{T}$, so that the goal is to predict the labels of the remaining nodes. For a semi-supervised node classification task on a graph, the labels $Y_{L}$ are only available for a subset of the noes (training set), so the goal is to predict the labels of the remaining nodes.
In GNN tasks, the topology information and node feature are combined to learn the representation vetor of a node for node tasks. Modern GNNs aggregate the information from nodes in neighborhood and update the representation of the nodes by a message-passing scheme. After k-iterations of aggregation, the representation of a node captures the structural information within its k-hop neighborhood. Formally, the layer-wise aggregation of a GNN is given by
$a_{v}^{(k)}=Aggregate^{(k)}(\{h_{u}^{k-1}:u\in N(v)\}),h_{v}^{(k)}=Combine^{(k)}(h_{v}^{(k-1)},a_{v}^{(k)})$ (1)
where $h_{v}^{(k)}$ is the feature vector of node v at the $k^{th}$ layer. We initialize the $h_{v}^{(0)}=x_{v}$, and the $\mathcal{N(v)}$ is a set of nodes connected to $v$. Existing work are focus on how to improve the work within 1-hop $Aggregate^{(k)}(\dot{)}$ and $Combine^{(k)}(\dot{)}$ operations. For example, GCN employs a convolution operation with the following rule
$H^{(l+1)}=\sigma(\hat{A}H^{(l)},\theta^{(l)})$ (2)
where $H^{(l)}$ are the output features in the $l^{th}$ layer. We initialize the $H^{(0)}=X$, $\hat{A}=\hat{D}^{-\frac{1}{2}}(A+I)\hat{D}^{-\frac{1}{2}}$ is the Laplacian normalized adjacency matrix, and the $\theta^{(l)}$ is the $l^{th}$ layer weights of the neural network. GNN models stack GNN layer to explore the representation in larger domain, take two-layer GCN model as an example, the represetnation can be calculated as $Z=softmax(\hat{A}\sigma(\hat{A}X\theta^{(0)})\theta^{(1)})$. As for node classification task with the labeled nodes corresponding to $Y_{L}$, the objective function is
$L=-\frac{1}{|\mathcal{Y_{L}}|}\sum_{v_{i}\in\mathcal{Y_{L}}}y_{i}log(z_{i})$ (3)
4.1 The In-Depth Learning of Graph Neural Networks
Contemporary Graph Neural Network (GNN) models predominantly adopt a 1-hop depth message passing paradigm, where each propagation cycle explores information at a singular depth level. In contrast to GraphSAGE [HYL17], which samples along k-depth but aggregates at every depth, our approach aims to develop a model capable of both sampling and aggregating in a k-hop domain. This method is anticipated to yield a more comprehensive understanding of the local structural context within the network.
We propose an in-depth learning algorithm for k-hop sampling and message passing in GNNs, characterized by the following steps:
In the proposed Graph Neural Network model, we initiate with a Static Preprocessing Step, where a substantial number of neighbors for each node are sampled. This phase involves constructing a k-hop adjacency matrix for every node, effectively capturing and storing the k-hop neighborhood information. This process is computationally efficient, with a complexity of $\mathcal{O}(\mathcal{V*K})$, where $\mathcal{V}$ represents the total number of vertices in the graph. Following this, the model enters the Dynamic Update Phase during propagation and update cycles. In this phase, a smaller, more focused subset of k-hop neighbors for each node is dynamically sampled. This targeted approach allows for the refinement of the pre-processed k-hop neighborhood matrix, ensuring that the model remains up-to-date and relevant in representing the evolving graph structure. The computational complexity of this phase is maintained at $\mathcal{O(K*R)}$, with $\mathcal{R}$ indicating the limited number of vertices sampled in the k-hop domain during each propagation step. Together, these steps create a robust framework for in-depth learning in Graph Neural Networks, optimizing both the breadth and depth of neighborhood information processing.
The k-hop attentive message passing strategy is underpinned by both dynamic and static sampling paths. This mechanism is grounded in the principles of graph diffusion networks, while the computation of distance reevaluation indicators is inspired by and derived from the architecture of the Graph Attention Network (GAT) [VCC+18]. This novel approach aims to harness the strengths of both static and dynamic sampling methodologies to enhance the depth and accuracy of message passing in GNNs.
Input :  Graph $\mathcal{G(V,E)}$; depth $K$; Sampling number per hop $N$; Input features: $Z$; Adjacency Matrix: A;
Output : NGH //K-hop Sampling Storage, a 3D dictionary
NGH[0][:] $\leftarrow$ A
// Static Sampling before Computation
for n in $V$ do
       for $k=K...1$ do
             $SN=RS(N,Layer^{(k-1)})$ for $v_{i}$ in layer(n,k-1) do
                   $NGH[k][n]\leftarrow NGH[k][n]\cup GraphSAGE(v_{i},SN[i])$
             end for
       end for
end for
// Dynamical Re-Sampling during Computation
$\mathcal{B^{K}}\leftarrow\mathcal{B}$; // The $\mathcal{B}$ denotes the current batch of nodes to be processed.for $k=K...1$ do
       $B^{k-1}\leftarrow B^{k}$; for $u_{i}\in\mathcal{B}$ do
             $SN_{s}\leftarrow RS(u,N)$$NGH[k][n]\leftarrow NGH[k][n]\cup GraphSAGE(u_{i},SN_{s}[i])$
            $z_{u_{i}}\leftarrow GKHDA(u_{i},k,NGH[k][n])$
       end for
end for
Algorithm 1 Preprocess Static Sampling and Dynamic Resampling Strategies
Input : Graph $\mathcal{G(V,E)}$;
input features: $x_{v}$
depth K;
non-linearity: $\sigma$;
differentiable aggregator functions $AGGREGATE_{k}$, $\forall k\in\{1,...,K\}$;
neighborhood sampling functions, $N_{k}:v\longrightarrow 2^{v}$, $\forall k\in\{1,...,K\}$;
NGH = $\leftarrow$ StaticSampling($\mathcal{G}$,K,Z,A);
weight matrices $W^{k}$; //learnable matrix
Output : Vector Representation $z_{v}$ for all $v\in\mathcal{B}$
$h_{u}^{0}\leftarrow x_{v}$, $\forall v\in B$ for $k=1...K$ do
       for $u\in B^{k}$ do
             $h^{k}_{u}\leftarrow\sigma(W^{k}\cdot(h_{u}^{k-1},h_{N(u)}^{k}))$; $h^{k}_{u}\leftarrow h^{k}_{u}/\lVert h^{k}_{u}\rVert_{2}$;
       end for
end for
for $k=K...1$ do
       for $u\in NGH[u,k]$ do
             $h^{k}_{N(u)}\leftarrow\mathcal{AGGREGATE}_{k}\{\mathcal{GAT}(h^{k-1}_{u^{\prime}},W^{k})\,\forall u^{\prime}\in N_{k}(u)\}$;
       end for
end for
$z_{u}\leftarrow\mathcal{COMBINE}(h_{u}^{i}),\forall i\in range(1,K+1)$
Algorithm 2 K-hop Diffusion Attention Layer Implementation
In our Graph Neural Network framework, the implementation of the $\mathcal{COMBINE}$ function plays a pivotal role. It can be expressed mathematically as:
$z_{u}\leftarrow W_{i}\cdot h_{u}^{i},\quad\forall i\in\text{range}(1,K+1)$ (4)
where $z_{u}$ represents the combined feature vector for a node $u$, and $h_{u}^{i}$ denotes the feature vector of the node at the $i$-th hop. $W_{i}$ is the weight matrix corresponding to the $i$-th hop, ensuring that features from different hops are weighted differently in the aggregation process.
Similarly, the $\mathcal{GAT}$ function, pivotal in our model for attention mechanism, is formulated as:
$\alpha_{ij}=\frac{\text{LeakyReLU}(\vec{a}^{T}[W\hat{h_{i}}\,||\,W\hat{h_{j}}])}{\sum_{k\in N_{i}}\text{LeakyReLU}(\vec{a}^{T}[W\hat{h_{i}}\,||\,W\hat{h_{k}}])}$ (5)
Here, $\alpha_{ij}$ is the attention coefficient between nodes $i$ and $j$, calculated using the LeakyReLU activation function. $\vec{a}$ is the attention vector and $W$ is the weight matrix applied to the feature vectors $\hat{h_{i}}$ and $\hat{h_{j}}$ of the nodes $i$ and $j$. The attention mechanism effectively captures the importance of each neighbor’s features in the aggregation process.
In this paper, we adopt a structure that considers both k-hop sampling and k-hop message passing within the graph diffusion network framework. Striking a balance between computational efficiency (time) and model performance, our approach involves static preprocessing for initial k-hop neighborhood sampling, followed by dynamic resampling in subsequent iterations. This methodology ensures that our model remains efficient while effectively capturing the complex dependencies in the graph structure across multiple hops.
4.2 Distance Recomputator and Topology Reconstructor Implementation
[…]
input features: $x_{v}$, $\forall v\in\mathcal{B}$;
depth $K$: weight matrices $W^{k}$
non-linearity: $\sigma$
differentiable aggregator functions $\text{AGGREGATE}_{k}$, $\forall k\in\{1,...,K\}$
neighborhood sampling functions, $N_{k}:v\longrightarrow 2^{v}$, $\forall k\in\{1,...,K\}$
recompute distance upper-bound: $\alpha$
recompute distance lower-bound: $\beta$
Sample factor: $\gamma$
Init Global NGH = $[0]_{V\times(k\times V)}$
Init Global W = $[\delta_{ij}]_{k\times F}$ //learnable matrix
Output : Vector Representation $z_{v}$ for all $v\in\mathcal{B}$
NGH = K-Hop-Sample($\mathcal{G}$)
for Each Propagation do
[…]
             DRM[i+1][:, :] $\leftarrow$ DRM[i+1][:, :] + DRM[i][DRM[i] ¿ $\alpha$]DRM[i][:, :] $\leftarrow$ DRM[i][:, :] - DRM[i][DRM[i] ¿ $\alpha$]
       end for
      for i = 1, …, K do
             DRM[i][:, :] $\leftarrow$ DRM[i][:, :] + DRM[i-1][DRM[i-1] ¡ $\beta$]DRM[i-1][:, :] $\leftarrow$ DRM[i-1][:, :] - DRM[i-1][DRM[i-1] ¡ $\beta$]
       end for
end for
Algorithm 3 Distance Recomputator and Asynchronous Aggregator Implementation
Distance Recomputator and Topology Reconstructor is meticulously designed to dynamically reconfigure graph topology and recalibrate node distances, thereby significantly enhancing the representational capacity and performance of GNNs.
Initial Setup and K-Hop Sampling: The algorithm begins with an initial setup phase where it prepares the graph $\mathcal{G(V,E)}$ with input features $x_{v}$ for all vertices $v$ in the batch $\mathcal{B}$. Each node is associated with a depth $K$, represented by weight matrices $W^{k}$, and a non-linearity function $\sigma$. The model incorporates differentiable aggregator functions $\mathcal{AGGREGATE}_{k}$ for each depth $k$, along with neighborhood sampling functions $N_{k}$. These components collectively form the foundation for k-hop neighborhood sampling, a crucial step in our model’s operation.
Dynamic Resampling and Recomputation: At the heart of our model lies a dynamic resampling process, which is executed in each propagation phase. This process adapts to the evolving graph structure by resampling a smaller subset of k-hop neighbors for each node. The resampled data, represented by NGH (Neighborhood Graph), is then refined through a series of computations to update the node representations effectively.
Distance Recomputator Mechanism: Central to our model is the Distance Recomputator (DRM), which recalculates the distances between nodes based on certain criteria, including recompute distance bounds $\mathcal{\alpha}$ and $\mathcal{\beta}$, and a sampling factor $\gamma$. This mechanism is crucial for adjusting the topological structure of the graph, ensuring that nodes are positioned optimally based on their relational context within the graph.
Graph Attention Network (GAT) Integration: The model integrates the Graph Attention Network (GAT) to compute attention coefficients between nodes. This integration allows the model to weigh the importance of each neighbor’s features in the aggregation process, further refining the distance recomputation and neighborhood sampling.
Asynchronous Aggregation: Another key aspect of our model is the implementation of an asynchronous aggregation approach. Unlike traditional methods that aggregate information simultaneously across all nodes, our model allows for selective and time-staggered aggregation based on the dynamic resampling and recomputation results. This approach ensures a more nuanced and efficient processing of graph data, crucial for handling large-scale and complex networks.
Application and Efficiency: The proposed algorithm is not only theoretically robust but also demonstrates practical efficiency in handling GNN tasks. By balancing the computational demands (time complexity) and model performance, our approach offers a feasible solution for real-world applications requiring in-depth graph analysis and dynamic topology reconstruction.
In summary, our algorithm presents a novel framework that combines k-hop sampling with dynamic resampling and recomputation, all underpinned by an asynchronous aggregation strategy. This comprehensive approach addresses key challenges in GNNs, such as handling complex graph topologies and efficiently updating node representations, thereby setting a new standard in the field of graph neural network research.
Figure 1: An instance of The Distance Computation and Topology Reconstruction
4.3 Experimental Analysis
Our experimental study focused on evaluating the performance of the proposed models - GKHDA, GDRA, and GKHDDRA - alongside typical scalable Graph Neural Network (GNN) methods, including GCN, SGC, S2GC, and APPNP. The experiments concentrated on graph structure learning and were conducted on benchmark datasets: Cora, Pubmed, and Citeseer.
Performance Overview: The results, as presented in Table LABEL:tab:widgets, showcase the efficacy of our models in comparison with established GNN methods like GAT, GraphSAGE, GCN, SGC, SSGC, and APPNP. Notably, our models - particularly when integrated with existing GNN structures (GCN, SGC, SSGC, and APPNP) - demonstrate superior performance across all datasets.
Dataset-Specific Analysis:
  • Cora: In the Cora dataset, the highest performance was observed with the APPNP+GKHDDRA combination, achieving an accuracy of 84.6%. This result indicates the robustness of GKHDDRA when combined with APPNP’s propagation scheme, which effectively leverages long-range dependencies in the graph.
  • Pubmed: The SGC+GKHDDRA combination outperformed other models with an accuracy of 82.5%. This suggests that the structured sparsity imposed by SGC, coupled with the advanced hop-wise learning capability of GKHDDRA, is particularly effective for the Pubmed dataset’s topology.
  • Citeseer: For Citeseer, the SSGC+GKHDDRA combination achieved the highest accuracy at 75.6%. This underscores the effectiveness of incorporating k-hop sampling and dynamic resampling in dealing with the dataset’s complex graph structure.
Comparative Assessment: The integration of our models with existing GNN frameworks consistently improved performance across datasets. For instance, GCN, when enhanced with GDRA and GKHDDRA, saw notable improvements in accuracy, emphasizing the value added by our distance recomputation and dynamic resampling mechanisms. Similarly, the integration with SGC and SSGC yielded significant performance boosts, highlighting the synergy between our models and scalable GNN methods in handling large-scale graph structures.
Model-Specific Contributions:
  • GKHDA showcased consistent improvements in graph representation learning, particularly in combination with SSGC and APPNP.
  • GDRA excelled in recalibrating the graph topology, which was evident from its strong performance, especially when combined with SGC and SSGC.
  • GKHDDRA emerged as a versatile model, enhancing both graph structure learning and node representation, as reflected in its top-tier results across all datasets, particularly when combined with APPNP.
Conclusion: The experimental results validate the effectiveness of our proposed models in enhancing the learning capabilities of GNNs. The integration of GKHDA, GDRA, and GKHDDRA with existing scalable GNN methods not only improved performance but also demonstrated their adaptability and compatibility with different graph structures and datasets. These findings indicate that our models are not only theoretically sound but also practically potent in a variety of real-world graph learning scenarios.
Figure 2: K-Hop SGC (Max Test Acc: 74.2%)
Figure 3: GDRA + K-Hop SGC (Max Test Acc: 75.8%)
Appendix A Appendix
Cora Pubmed Citeseer
GAT 82.1% 77.7% 69%
GraphSAGE 81.5% 70.3% 74%
GDRA 82.5% 77.8% 69.8%
Table 1: Comparision of GDRA and its basic two components (GAT and GraphSAGE)
Cora Pubmed Citeseer
GKHDA 82.3% 78.6% 70.7%
GDRA 82.5% 77.8% 69.8%
GKHDDRA 82.4% 79.4% 71.2%
Table 2: Comparison of 3 Basic Model of DRTR
Cora Pubmed Citeseer
GCN 81.2% 79.3% 70.9%
GCN+GDRA 82.6% 80.1% 71.3%
GCN+GKHDA 82.4% 80.5% 71.7%
GCN+GKHDDRA 82.7% 80.9% 72.3%
Table 3: GCN and SGC + DRTR + Diff comparison
Cora Pubmed Citeseer
SGC 74.2% 78.2% 71.5%
SGC+GDRA 75.8% 81.2% 73.1%
SGC+GKHDA 75.1% 81.6% 73.4%
SGC+GKHDDRA 77.4% 82.5% 74.6%
Table 4: SGC and SGC + DRTR + Diff comparison
Cora Pubmed Citeseer
SSGC 83.0% 73.6% 75.6%
SSGC+GDRA 83.2% 74.2% 76.4%
SSGC+GKHDA 84.3% 74.5% 76.1%
SSGC+GKHDDRA 84.1% 74.7% 77.6%
Table 5: SSGC and SSGC + DRTR + Diff comparison
Cora Pubmed Citeseer
APPNNP 82.3% 71.5% 75.2%
APPNP+GDRA 83.5% 73.6% 74.4%
APPNP+GKHDA 83.8% 74.1% 74.5%
APPNP+GKHDDRA 84.6% 74.5% 75.3%
Table 6: APPNP and APPNP + DRTR + Diff comparison
Dataset Nodes Edges Features Classes
PubMed 19,717 44,338 500 3
Cora 2,708 5,429 1,433 7
CiteSeer 3,312 4,732 3,703 6
Table 7: Comparison of PubMed, Cora, and CiteSeer in Terms of Nodes, Edges, Features, and Classes
Models DR TR k_hop_resampling Heat_Diffusion_Propagation
GDRA $\checkmark$ $\checkmark$
GKHDA $\checkmark$ $\checkmark$
GKHDDRA $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 8: GDRA, GKHDRA, GKHDDRA Modules (DR: Distance Recomputator; TR: Topology Reconstructor)
Table 9: Experimental Settings
Learning Rate Weight Decay Epochs Patience
0.005 0.001 1000 100