### v3

# Adaptive Topology Reconstruction for Robust Graph Representation Learning

Abstract
Graph Neural Networks (GNNs) have become fundamental in semi-supervised learning for graph representation, leveraging their ability to capture complex node relationships. A recent trend in GNN research focuses on adaptive multi-hop structure learning, moving beyond fixed-hop aggregation to more flexible and dynamic neighborhood selection. While GAMLP (Zhang et al., 2022) employs separate MLP layers for each multi-hop domain and ImprovingTE (Yao et al., 2023) enhances this by injecting contextualized substructure information, these methods still rely heavily on predefined sampling strategies, which may limit their ability to generalize and maintain stable accuracy. To address these limitations, we propose an adaptive reconstruction framework that dynamically refines multi-hop structure learning. Inspired by ”coreset selection” (Guo et al., 2022), our approach adaptively reconstructs node neighborhoods to optimize message passing, ensuring more effective and context-aware information flow across the graph. To further enhance structural robustness, we introduce two key modules: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator reassesses and recalibrates node distances based on adaptive graph properties, leading to improved node embeddings that better reflect latent relationships. Meanwhile, the Topology Reconstructor dynamically refines local graph structures, enabling the model to adapt to evolving graph topologies and mitigate the impact of noise and mislabeled data. Empirical evaluations demonstrate that our adaptive reconstruction framework achieves significant improvements over existing multi-hop-based models, providing more stable and accurate performance in various graph learning benchmarks.
1 Introduction
Existing GNN architectures predominantly adopt static node representations and predefined aggregation schemes, which can be suboptimal when handling graphs with dynamic topologies. Furthermore, the computation of node distances in large-scale graphs introduces significant overhead, making it challenging to scale these models effectively (Kung et al., 2024; Wang et al., 2024). These limitations highlight the need for a more flexible and efficient approach that can dynamically reconstruct graph structures, allowing models to better adapt to changes in connectivity patterns and improve representation quality.
To address these challenges, we propose an adaptive reconstruction framework that refines both node distance calculations and local topological structures to enhance message passing efficiency. This framework introduces two key components: the Distance Recomputator and the Topology Reconstructor (DRTR). The Distance Recomputator dynamically recalibrates node distances within the multi-hop domain by incorporating an adaptive encoding mechanism that adjusts to changes in graph density and node proximity. By refining distance computations, it provides a more precise representation of node relationships, which improves downstream predictions and model stability. The Topology Reconstructor complements this by continuously adjusting local graph connectivity, ensuring that the model remains robust in scenarios where structural variations occur. This adaptive mechanism allows GNNs to dynamically reconstruct neighborhood structures, making them more resilient to graph shifts and variations in node interactions.
Through extensive experiments on multiple benchmark datasets, we demonstrate that our adaptive reconstruction approach significantly improves performance in tasks involving both static and dynamic graphs. The Distance Recomputator proves effective in optimizing node representations, reducing the impact of noisy or misleading edges while enhancing the quality of learned embeddings. Meanwhile, the Topology Reconstructor enables models to better align with evolving graph structures, leading to more stable and accurate predictions. In addition to empirical validation, we provide a theoretical analysis that explains the benefits of reconstruction in GNN learning dynamics, offering deeper insights into why adaptively refining graph structures leads to superior model generalization.
2 Motivation
2.1 Topology Imbalance Problem
The structural connections in a graph do not always accurately reflect the significance of each link. As stated in (Liu et al., 2023), ”not every link is as useful as its topology suggests.” Consider a social network where an account labeled ”math” follows ten mathematical educators and one food-maker. For tasks such as label prediction, the connection to the food-maker contributes little relevant information, yet traditional Graph Neural Networks (GNNs) aggregate information from all 1-hop neighbors indiscriminately. This results in a phenomenon we define as the Topology Imbalance Problem, where some connections (e.g., links to math educators) carry significantly more useful information than others (e.g., the food-maker link).
While models such as Graph Attention Networks (GAT) assign importance scores to each link, they still rely on synchronous aggregation, where all nodes within the same hop are processed simultaneously. This limitation prevents the model from prioritizing more relevant connections over less meaningful ones during message passing. To overcome this issue, we propose the Distance Recomputator, which dynamically recalculates node distances to better reflect the contextual importance of each link. Additionally, we introduce the Asynchronous Aggregator, which enables nodes to be aggregated based on these recomputed distances, allowing for a more adaptive and information-sensitive update mechanism.
2.2 Limitations of Synchronous Aggregation
Most GNN architectures employ synchronous aggregators, where information from a node’s immediate neighbors is aggregated simultaneously. For example, Graph Convolutional Networks (GCN) aggregate neighbor information by weighting it with the inverse node degree, while GAT uses an adjacency matrix mask and attention scores to prioritize neighbors. Similarly, GraphSAGE samples a subset of neighbors before applying an aggregation function. Despite their differences in implementation, these models share a common characteristic: all nodes in the same layer are processed concurrently.
A major drawback of synchronous aggregation is its inability to differentiate between links of varying importance in a timely manner. Since all neighboring nodes are updated simultaneously, irrelevant or even erroneous connections can propagate misleading information, reducing model robustness. This issue becomes more pronounced in large and complex graphs, where spurious or noisy edges can significantly degrade performance.
To address this limitation, we propose an asynchronous aggregation mechanism that selectively integrates information from different neighbors based on their relevance. Unlike traditional synchronous approaches, our method dynamically determines the order and priority of aggregation, reducing the negative influence of uninformative links while preserving critical structural relationships. Some recent works have explored asynchronous processing in GNNs, such as AEGNN (Schaefer et al., 2022), which updates node activations only upon new events, and Gated Graph Sequence Neural Networks (Li et al., 2017), which employ gated recurrent units to process sequences asynchronously. However, these methods primarily focus on message passing rather than addressing topological reconstruction at a structural level. Our approach directly enhances the representation of graph topology, offering a more comprehensive solution.
2.3 Heat Diffusion in Graph Neural Networks
The concept of heat diffusion offers a valuable perspective for understanding information propagation in Graph Neural Networks (GNNs). Drawing inspiration from thermodynamics, where heat gradually dissipates as it moves away from its source (Thanou et al., 2016), we analyze how information in GNNs follows a similar pattern of attenuation during multi-hop message passing. As information propagates across a graph, its influence weakens with increasing distance from the source node, often leading to a decline in signal quality and an accumulation of noise. Traditional GNN architectures typically assume uniform propagation across all hops, failing to account for this natural decay in information relevance.
To address this limitation, we introduce a distance-aware reconstruction framework that models message passing through a heat diffusion-inspired mechanism. Our approach regulates the transmission strength based on a distance-dependent attenuation function, ensuring that information decays in a controlled and meaningful manner. By integrating this adaptive reconstruction process, we prevent the excessive spread of weak or irrelevant signals while reinforcing the retention of important structural patterns. This refinement enhances both the signal-to-noise ratio and the interpretability of learned node representations, making the model more robust to graph sparsity and noise.
By incorporating heat diffusion principles into the reconstruction process, our method achieves a more balanced and adaptive message propagation strategy. This approach not only improves the efficiency of information flow but also mitigates common issues such as oversmoothing and excessive feature mixing in deep GNN models. Additional theoretical analysis and mathematical formulations of our heat diffusion-based reconstruction mechanism are provided in our supplementary document, ”Heat Diffusion in GNNs.”
2.4 Paper Contribution
  • K-hop Diffusion Message Passing: We propose a message passing framework that integrates k-hop information at each propagation step. After each update, the model recalibrates node embeddings and distances to reflect newly acquired information. A heat diffusion-inspired mechanism controls attenuation, preserving meaningful connections while reducing noise.
  • Distance Recomputator: Our model recalculates node distances using an adaptive attention mechanism that considers both multi-hop topology and feature interactions. It continuously adjusts distances based on contextual relevance, improving message passing and filtering irrelevant edges.
  • Topology Reconstructor: We introduce a multi-hop topology reconstruction method. our model leverages multi-hop topology from sampling to perform reconstruction based on computed ”similarity distances.” Nodes exceeding a similarity distance threshold are repositioned to optimize network configurations for enhanced learning outcomes.
3 Core Ideas of DRTR
3.1 K-hop Message Passing
Most existing multi-hop models in graph neural networks rely on multiple rounds of 1-hop message passing to capture higher-order information. However, they lack the ability to perform k-hop message passing with multi-hop propagation in a single iteration or efficiently sample multi-hop neighborhoods in one step. For example, the enhancing multi-hop connectivity model refines multi-hop neighborhoods by sampling multi-hop neighbors and reweighting high-order connections, strengthening highly related nodes while down-weighting weaker ones (Liu et al., 2022). Similarly, KP-GNN enriches node representations by incorporating peripheral embeddings at each layer (Feng et al., 2022).
Our model introduces an efficient multi-hop sampling strategy that enables direct k-hop message passing, capturing local structures more effectively than traditional 1-hop methods. By integrating multi-hop information in each propagation step, our approach provides a more comprehensive graph perspective, leading to improved label prediction accuracy for central nodes.
3.2 Diffusion Propagation in Graph Neural Networks
We propose a diffusion-based propagation model that builds upon traditional approaches such as the Diffusion Decent Network. One well-known model, MAGNA (Wang et al., 2021), propagates information by updating vertex and edge embeddings while recalculating distances for both one-hop and multi-hop neighbors, considering all paths up to $k$ hops. Another approach, ”Diffusion Improves Graph Learning” (Klicpera et al., 2019), approximates the diffusion equation using an infinite series, improving computational efficiency on large-scale graphs.
Our model extends these concepts by simulating heat diffusion within the multi-hop domain. This diffusion-inspired framework regulates message passing, dynamically adjusting the flow of information to reduce noise accumulation and enhance representation quality. By incorporating heat diffusion into multi-hop propagation, our model provides a structured and adaptive approach to information distribution in graph neural networks.
3.3 Graph Imbalance Problems
Label and topology imbalance in graph neural networks remain significant but underexplored challenges. Zhao et al. (Zhao et al., 2022) addressed topology imbalance by adjusting edge weights without modifying the graph structure.
We introduce two models that actively reconstruct graph topology to improve balance. The first, GKHDDRA, applies hop jumping to restructure the graph and refine connectivity. The second, GDRA, enhances the dataset by removing weak edges and forming new connections between strongly similar nodes. These adjustments optimize graph structure, leading to a more balanced and representative network, ultimately improving learning performance.
4 Architectures of DRTR
The Distance Recomputator and Topology Reconstructor (DRTR) is designed to enhance Graph Neural Networks (GNNs) by addressing two major challenges: (1) capturing long-range dependencies while maintaining computational efficiency, and (2) dynamically adjusting the graph topology based on learned representations. The K-hop Diffusion Attention Layer, a core component of DRTR, integrates static preprocessing, multi-hop diffusion, and dynamic topology updates to optimize message passing and representation learning.
Graph-based learning problems often rely on an underlying undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, with the node set $\mathcal{V}$ and edge set $\mathcal{E}$. The adjancy matrix A $\in$ $R^{N\times N}$ describes the existence of connections among nodes. Every node $v_{i}\in\mathcal{V}$ has an associated feature vector $x_{i}\in R^{d\times 1}$, so the whole feature vector space can be represented as $X=[x_{1},x_{2},...,x_{N}]^{T}$, so that the goal is to predict the labels of the remaining nodes. For a semi-supervised node classification task on a graph, the labels $Y_{L}$ are only available for a subset of the noes (training set), so the goal is to predict the labels of the remaining nodes.
In GNN tasks, the topology information and node feature are combined to learn the representation vetor of a node for node tasks. Modern GNNs aggregate the information from nodes in neighborhood and update the representation of the nodes by a message-passing scheme. After k-iterations of aggregation, the representation of a node captures the structural information within its multi-hop neighborhood. Formally, the layer-wise aggregation of a GNN is given by
$a_{v}^{(k)}=Aggregate^{(k)}(\{h_{u}^{k-1}:u\in N(v)\}),h_{v}^{(k)}=Combine^{(k)}(h_{v}^{(k-1)},a_{v}^{(k)})$ (1)
where $h_{v}^{(k)}$ represents the feature vector of node $v$ at the $k^{th}$ layer. The initialization follows $h_{v}^{(0)}=x_{v}$, where $\mathcal{N}(v)$ denotes the set of neighboring nodes connected to $v$.
Existing works primarily focus on improving the 1-hop aggregation $\mathcal{AGGREGATE}^{(k)}(\cdot)$ and combination $\mathcal{COMBINE}^{(k)}(\cdot)$ operations. For example, Graph Convolutional Networks (GCN) utilize a convolution operation defined as:
$H^{(l+1)}=\sigma(\hat{A}H^{(l)}\Theta^{(l)})$ (2)
where $H^{(l)}$ represents the feature matrix at the $l^{th}$ layer, initialized as $H^{(0)}=X$. The term $\hat{A}=\hat{D}^{-\frac{1}{2}}(A+I)\hat{D}^{-\frac{1}{2}}$ is the Laplacian-normalized adjacency matrix, and $\Theta^{(l)}$ is the learnable weight matrix at the $l^{th}$ layer.
GNN models stack multiple layers to capture hierarchical representations over a larger neighborhood. Taking a two-layer GCN as an example, the final node representation is computed as:
$Z=\text{softmax}(\hat{A}\sigma(\hat{A}X\Theta^{(0)})\Theta^{(1)})$ (3)
For a node classification task, given labeled nodes corresponding to $Y_{L}$, the objective function is defined as:
$L=-\frac{1}{|\mathcal{Y_{L}}|}\sum_{v_{i}\in\mathcal{Y_{L}}}y_{i}log(z_{i})$ (4)
4.1 In-Depth Learning in Graph Neural Networks
Contemporary GNNs predominantly employ a 1-hop message passing paradigm, where each propagation step explores information within a single depth level. Unlike GraphSAGE (Hamilton et al., 2017), which samples along a $k$-depth domain but aggregates at every depth, our approach integrates both sampling and aggregation within a $k$-hop neighborhood. This enables a more comprehensive understanding of local structural dependencies within the graph.
We propose an in-depth learning algorithm for $k$-hop sampling and message passing in GNNs, comprising the following key components:
1. Static Preprocessing Step: A large number of neighbors are pre-sampled for each node, constructing a $k$-hop adjacency matrix that encapsulates extended neighborhood information. This process achieves efficient storage of structural dependencies with a computational complexity of $\mathcal{O}(\mathcal{V}K)$, where $\mathcal{V}$ denotes the total number of vertices.
2. Dynamic Update Phase: During propagation, a refined subset of $k$-hop neighbors is dynamically sampled, updating the precomputed neighborhood structure. This step ensures that the model adapts to evolving graph topologies while maintaining computational efficiency at $\mathcal{O}(KR)$, where $\mathcal{R}$ represents the number of sampled vertices per propagation step.
The $k$-hop attentive message passing strategy leverages both static and dynamic sampling pathways. This mechanism is inspired by graph diffusion networks and integrates distance reevaluation indicators derived from the Graph Attention Network (GAT) (Veličković et al., 2018). By synthesizing static and dynamic sampling methodologies, our approach enhances both the depth and accuracy of message passing in GNNs.
Input :  Graph $\mathcal{G(V,E)}$, Depth $K$, Sampling number per hop $N$,
Input features $\mathbf{Z}$, Adjacency Matrix $\mathbf{A}$
Output : $\mathcal{NGH}$: K-hop Sampling Storage (3D dictionary)
$\mathcal{NGH}^{(0)}\leftarrow\mathbf{A}$
  // Initialize neighborhood
// Static Sampling before Computation
for $n\in\mathcal{V}$ do
       for $k=K\dots 1$ do
             $\mathcal{SN}=\mathcal{RS}(N,\text{Layer}^{(k-1)})$
             for $v_{i}\in\text{Layer}(n,k-1)$ do
                   $\mathcal{NGH}^{(k)}_{n}\leftarrow\mathcal{NGH}^{(k)}_{n}\cup\text{GraphSAGE}(v_{i},\mathcal{SN}[i])$
             end for
       end for
end for
// Dynamical Re-Sampling during Computation
$\mathcal{B}^{(K)}\leftarrow\mathcal{B}$ // $\mathcal{B}$: Batch of nodes to process
for $k=K\dots 1$ do
       $\mathcal{B}^{(k-1)}\leftarrow\mathcal{B}^{(k)}$
       for $u_{i}\in\mathcal{B}$ do
             $\mathcal{SN}_{s}\leftarrow\mathcal{RS}(u_{i},N)$
             $\mathcal{NGH}^{(k)}_{n}\leftarrow\mathcal{NGH}^{(k)}_{n}\cup\text{GraphSAGE}(u_{i},\mathcal{SN}_{s}[i])$
             $z_{u_{i}}\leftarrow\mathcal{GKHDA}(u_{i},k,\mathcal{NGH}^{(k)}_{n})$
       end for
end for
Algorithm 1 Static Sampling and Dynamic Resampling
DRTR Diffusion Layer design is an integration of multi-hop diffusion, adaptive message passing, and dynamic topology reconstruction to capture both local and long-range dependencies. Instead of standard 1-hop aggregation, DRTR employs an iterative K-hop aggregation process, where node representations are refined using a combination of local neighborhood aggregation and global diffusion attention. To further improve feature propagation, a Graph Attention Mechanism (GAT) dynamically computes attention weights, prioritizing important neighbors and filtering out noisy connections. Additionally, DRTR introduces a Topology Reconstructor Module, which adjusts edge weights and graph connectivity by recomputing distances and adapting to evolving structures. This mechanism mitigates over-smoothing and ensures the model remains responsive to structural variations. By balancing computational efficiency and expressive power, DRTR’s hierarchical multi-hop framework enables effective large-scale graph learning while maintaining a trade-off between depth (global context) and locality (fine-grained details).
Input :  Graph $\mathcal{G(V,E)}$, Depth $K$, Non-linearity $\sigma$,
Input features $\mathbf{x_{v}}$, Weight matrices $\mathbf{W^{k}}$,
Aggregator functions $\mathcal{AGGREGATE}_{k}$, $\forall k\in\{1,...,K\}$,
Neighborhood sampling $N_{k}:v\to 2^{v}$, $\forall k\in\{1,...,K\}$,
$\mathcal{NGH}\leftarrow\text{StaticSampling}(\mathcal{G},K,\mathbf{Z},\mathbf{A})$
Output : Vector Representation $\mathbf{z_{v}}$ for all $v\in\mathcal{B}$
// Static Aggregation
$h_{u}^{0}\leftarrow x_{v}$, $\forall v\in\mathcal{B}$
for $k=1\dots K$ do
       for $u\in\mathcal{B}^{(k)}$ do
             $h^{k}_{u}\leftarrow\sigma(\mathbf{W}^{k}\cdot(h_{u}^{k-1},h_{N(u)}^{k}))$
             $h^{k}_{u}\leftarrow\frac{h^{k}_{u}}{\lVert h^{k}_{u}\rVert_{2}}$
       end for
end for
// K-hop Diffusion Attention
for $k=K\dots 1$ do
       for $u\in\mathcal{NGH}^{(k)}_{u}$ do
             $h^{k}_{N(u)}\leftarrow\mathcal{AGGREGATE}_{k}\{\mathcal{GAT}(h^{k-1}_{u^{\prime}},\mathbf{W}^{k})\,|\,u^{\prime}\in N_{k}(u)\}$
       end for
end for
$\mathbf{z_{u}}\leftarrow\mathcal{COMBINE}(h_{u}^{i}),\forall i\in\{1,\dots,K\}$
Algorithm 2 K-hop Diffusion Attention Layer Implementation
In our Graph Neural Network framework, the implementation of the $\mathcal{COMBINE}$ function plays a pivotal role. It can be expressed mathematically as:
$z_{u}\leftarrow W_{i}\cdot h_{u}^{i},\quad\forall i\in\text{range}(1,K+1)$ (5)
where $z_{u}$ represents the combined feature vector for a node $u$, and $h_{u}^{i}$ denotes the feature vector of the node at the $i$-th hop. $W_{i}$ is the weight matrix corresponding to the $i$-th hop, ensuring that features from different hops are weighted differently in the aggregation process.
Similarly, the $\mathcal{GAT}$ function, pivotal in our model for attention mechanism, is formulated as:
$\alpha_{ij}=\frac{\text{LeakyReLU}(\vec{a}^{T}[W\hat{h_{i}}\,||\,W\hat{h_{j}}])}{\sum_{k\in N_{i}}\text{LeakyReLU}(\vec{a}^{T}[W\hat{h_{i}}\,||\,W\hat{h_{k}}])}$ (6)
Here, $\alpha_{ij}$ is the attention coefficient between nodes $i$ and $j$, calculated using the LeakyReLU activation function. $\vec{a}$ is the attention vector and $W$ is the weight matrix applied to the feature vectors $\hat{h_{i}}$ and $\hat{h_{j}}$ of the nodes $i$ and $j$. The attention mechanism effectively captures the importance of each neighbor’s features in the aggregation process.
In this paper, we adopt a structure that considers both k-hop sampling and k-hop message passing within the graph diffusion network framework. Striking a balance between computational efficiency (time) and model performance, our approach involves static preprocessing for initial multi-hop neighborhood sampling, followed by dynamic resampling in subsequent iterations. This methodology ensures that our model remains efficient while effectively capturing the complex dependencies in the graph structure across multiple hops.
4.2 Distance Recomputator and Topology Reconstructor Implementation
Distance Recomputator and Topology Reconstructor is meticulously designed to dynamically reconfigure graph topology and recalibrate node distances, thereby significantly enhancing the representational capacity and performance of GNNs.
Input :  Graph $\mathcal{G(V,E)}$, Depth $K$, Non-linearity $\sigma$,
Input features $\mathbf{x_{v}}$, Weight matrices $\mathbf{W^{k}}$,
Aggregator functions $\mathcal{AGGREGATE}_{k}$, $\forall k\in\{1,...,K\}$,
Neighborhood sampling $N_{k}:v\to 2^{v}$, $\forall k\in\{1,...,K\}$,
Distance bounds $\alpha,\beta$, Sampling factor $\gamma$,
$\mathcal{NGH}\leftarrow[0]_{V\times(K\times V)}$, $\mathbf{W}\leftarrow[\delta_{ij}]_{K\times F}$
Output : Vector Representation $\mathbf{z_{v}}$ for all $v\in\mathcal{B}$
// K-hop Sampling
$\mathcal{NGH}\leftarrow\text{K-Hop-Sample}(\mathcal{G})$
for Each Propagation do
       $\mathcal{NGH}\leftarrow\mathcal{NGH}\cup\text{Resample}(\mathcal{G},\text{ResampleNum})$
       $\mathcal{NGH}_{\text{compute}}\leftarrow\text{Sample}(\mathcal{NGH},\gamma)$
       $\mathbf{Z_{v}}\leftarrow\text{GKHDDRA}(\mathcal{G},\mathcal{NGH}_{\text{compute}})$
       $\mathbf{DRM}\leftarrow\text{mask}(\mathcal{GAT}(\mathcal{G},\mathbf{Z_{v}}),\mathcal{NGH})$
      // Distance Recomputing
       for $i=K-1\dots 0$ do
             $\mathbf{DRM}[i+1]\leftarrow\mathbf{DRM}[i+1]+\mathbf{DRM}[i][\mathbf{DRM}[i]>\alpha]$
             $\mathbf{DRM}[i]\leftarrow\mathbf{DRM}[i]-\mathbf{DRM}[i][\mathbf{DRM}[i]>\alpha]$
       end for
      // Topology Reconstruction
       for $i=1\dots K$ do
             $\mathbf{DRM}[i]\leftarrow\mathbf{DRM}[i]+\mathbf{DRM}[i-1][\mathbf{DRM}[i-1]<\beta]$
             $\mathbf{DRM}[i-1]\leftarrow\mathbf{DRM}[i-1]-\mathbf{DRM}[i-1][\mathbf{DRM}[i-1]<\beta]$
       end for
end for
Algorithm 3 Distance Recomputator and Topology Reconstructor
Initial Setup and K-Hop Sampling: The algorithm begins with an initial setup phase where it prepares the graph $\mathcal{G(V,E)}$ with input features $x_{v}$ for all vertices $v$ in the batch $\mathcal{B}$. Each node is associated with a depth $K$, represented by weight matrices $W^{k}$, and a non-linearity function $\sigma$. The model incorporates differentiable aggregator functions $\mathcal{AGGREGATE}_{k}$ for each depth $k$, along with neighborhood sampling functions $N_{k}$. These components collectively form the foundation for k-hop neighborhood sampling, a crucial step in our model’s operation.
Dynamic Resampling and Recomputation: At the heart of our model lies a dynamic resampling process, which is executed in each propagation phase. This process adapts to the evolving graph structure by resampling a smaller subset of k-hop neighbors for each node. The resampled data, represented by NGH (Neighborhood Graph), is then refined through a series of computations to update the node representations effectively.
Distance Recomputator Mechanism: Central to our model is the Distance Recomputator (DRM), which recalculates the distances between nodes based on certain criteria, including recompute distance bounds $\mathcal{\alpha}$ and $\mathcal{\beta}$, and a sampling factor $\gamma$. This mechanism is crucial for adjusting the topological structure of the graph, ensuring that nodes are positioned optimally based on their relational context within the graph.
Graph Attention Network (GAT) Integration: The model integrates the Graph Attention Network (GAT) to compute attention coefficients between nodes. This integration allows the model to weigh the importance of each neighbor’s features in the aggregation process, further refining the distance recomputation and neighborhood sampling.
Asynchronous Aggregation: Another key aspect of our model is the implementation of an asynchronous aggregation approach. Unlike traditional methods that aggregate information simultaneously across all nodes, our model allows for selective and time-staggered aggregation based on the dynamic resampling and recomputation results. This approach ensures a more nuanced and efficient processing of graph data, crucial for handling large-scale and complex networks.
In summary, our algorithm presents a novel framework that combines multi-hop sampling with dynamic resampling and recomputation, all underpinned by an asynchronous aggregation strategy. This comprehensive approach addresses key challenges in GNNs, such as handling complex graph topologies and efficiently updating node representations, thereby setting a new standard in the field of graph neural network research.
Figure 1: An instance of The Distance Computation and Topology Reconstruction
4.3 Experimental Analysis
Our experimental study focused on evaluating the performance of the proposed models - GKHDA, GDRA, and GKHDDRA - alongside typical scalable Graph Neural Network (GNN) methods, including GCN, SGC, S2GC, and APPNP. The experiments concentrated on graph structure learning and were conducted on benchmark datasets: Cora, Pubmed, and Citeseer.
Performance Overview: The results, as presented in Table 2, showcase the efficacy of our models in comparison with established GNN methods like GAT, GraphSAGE, GCN, SGC, SSGC, and APPNP. Notably, our models - particularly when integrated with existing GNN structures (GCN, SGC, SSGC, and APPNP) - demonstrate superior performance across all datasets.
Dataset-Specific Analysis:
  • Cora: In the Cora dataset, the highest performance was observed with the APPNP+GKHDDRA combination, achieving an accuracy of 84.6%. This result indicates the robustness of GKHDDRA when combined with APPNP’s propagation scheme, which effectively leverages long-range dependencies in the graph.
  • Pubmed: The SGC+GKHDDRA combination outperformed other models with an accuracy of 82.5%. This suggests that the structured sparsity imposed by SGC, coupled with the advanced hop-wise learning capability of GKHDDRA, is particularly effective for the Pubmed dataset’s topology.
  • Citeseer: For Citeseer, the SSGC+GKHDDRA combination achieved the highest accuracy at 75.6%. This underscores the effectiveness of incorporating multi-hop sampling and dynamic resampling in dealing with the dataset’s complex graph structure.
Comparative Assessment: The integration of our models with existing GNN frameworks consistently improved performance across datasets. For instance, GCN, when enhanced with GDRA and GKHDDRA, saw notable improvements in accuracy, emphasizing the value added by our distance recomputation and dynamic resampling mechanisms. Similarly, the integration with SGC and SSGC yielded significant performance boosts, highlighting the synergy between our models and scalable GNN methods in handling large-scale graph structures.
Model-Specific Contributions:
  • GKHDA showcased consistent improvements in graph representation learning, particularly in combination with SSGC and APPNP.
  • GDRA excelled in recalibrating the graph topology, which was evident from its strong performance, especially when combined with SGC and SSGC.
  • GKHDDRA emerged as a versatile model, enhancing both graph structure learning and node representation, as reflected in its top-tier results across all datasets, particularly when combined with APPNP.
Conclusion: The experimental results validate the effectiveness of our proposed models in enhancing the learning capabilities of GNNs. The integration of GKHDA, GDRA, and GKHDDRA with existing scalable GNN methods not only improved performance but also demonstrated their adaptability and compatibility with different graph structures and datasets. These findings indicate that our models are not only theoretically sound but also practically potent in a variety of real-world graph learning scenarios.
Figure 2: K-Hop SGC (Max Test Acc: 74.2%)
Figure 3: GDRA + K-Hop SGC (Max Test Acc: 75.8%)
Appendix A Appendix
Cora Pubmed Citeseer
GAT 82.1% 77.7% 69%
GraphSAGE 81.5% 70.3% 74%
GDRA 82.5% 77.8% 69.8%
Table 1: Comparision of GDRA and its basic two components (GAT and GraphSAGE)
Cora Pubmed Citeseer
GKHDA 82.3% 78.6% 70.7%
GDRA 82.5% 77.8% 69.8%
GKHDDRA 82.4% 79.4% 71.2%
Table 2: Comparison of 3 Basic Model of DRTR
Cora Pubmed Citeseer
GCN 81.2% 79.3% 70.9%
GCN+GDRA 82.6% 80.1% 71.3%
GCN+GKHDA 82.4% 80.5% 71.7%
GCN+GKHDDRA 82.7% 80.9% 72.3%
Table 3: GCN and SGC + DRTR + Diff comparison
Cora Pubmed Citeseer
SGC 74.2% 78.2% 71.5%
SGC+GDRA 75.8% 81.2% 73.1%
SGC+GKHDA 75.1% 81.6% 73.4%
SGC+GKHDDRA 77.4% 82.5% 74.6%
Table 4: SGC and SGC + DRTR + Diff comparison
Cora Pubmed Citeseer
SSGC 83.0% 73.6% 75.6%
SSGC+GDRA 83.2% 74.2% 76.4%
SSGC+GKHDA 84.3% 74.5% 76.1%
SSGC+GKHDDRA 84.1% 74.7% 77.6%
Table 5: SSGC and SSGC + DRTR + Diff comparison
Cora Pubmed Citeseer
APPNNP 82.3% 71.5% 75.2%
APPNP+GDRA 83.5% 73.6% 74.4%
APPNP+GKHDA 83.8% 74.1% 74.5%
APPNP+GKHDDRA 84.6% 74.5% 75.3%
Table 6: APPNP and APPNP + DRTR + Diff comparison
Dataset Nodes Edges Features Classes
PubMed 19,717 44,338 500 3
Cora 2,708 5,429 1,433 7
CiteSeer 3,312 4,732 3,703 6
Table 7: Comparison of PubMed, Cora, and CiteSeer in Terms of Nodes, Edges, Features, and Classes
Models DR TR k_hop_resampling Heat_Diffusion_Propagation
GDRA $\checkmark$ $\checkmark$
GKHDA $\checkmark$ $\checkmark$
GKHDDRA $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 8: GDRA, GKHDRA, GKHDDRA Modules (DR: Distance Recomputator; TR: Topology Reconstructor)
Table 9: Experimental Settings
Learning Rate Weight Decay Epochs Patience
0.005 0.001 1000 100